# State-Covering Trajectory Stitching for Diffusion Planners

### v2

@@ Abstract
Diffusion-based generative models are emerging as powerful tools for long-horizon planning in reinforcement learning (RL), particularly with offline datasets. However, their performance is fundamentally limited by the quality and diversity of training data. This often restricts their generalization to tasks outside their training distribution or longer planning horizons. To overcome this challenge, we propose State-Covering Trajectory Stitching (SCoTS), a novel reward-free trajectory augmentation method that incrementally stitches together short trajectory segments, systematically generating diverse and extended trajectories. SCoTS first learns a temporal distance-preserving latent representation that captures the underlying temporal structure of the environment, then iteratively stitches trajectory segments guided by directional exploration and novelty to effectively cover and expand this latent space. We demonstrate that SCoTS significantly improves the performance and generalization capabilities of diffusion planners on offline goal-conditioned benchmarks requiring stitching and long-horizon reasoning. Furthermore, augmented trajectories generated by SCoTS significantly improve the performance of widely used offline goal-conditioned RL algorithms across diverse environments.
@@ 1 Introduction
In many real-world applications, agents must plan over hundreds of steps, often receiving sparse or delayed feedback until they reach a distant goal. Perfect knowledge of the environment allows powerful planners like MPC Tassa et al. (2012) and MCTS Silver et al. (2016, 2017); Lee et al. (2018) to excel. However, most real-world tasks instead require learning environment dynamics from data. Model-based reinforcement learning (MBRL) Sutton (2018) constructs such world models, offering sample-efficient learning and improved generalization Ha & Schmidhuber (2018); Hafner et al. (2019); Kaiser et al. (2020). However, autoregressive predictions from learned models accumulate small errors into a cascade of inaccuracies. This compounding error can cause planners to exploit model inaccuracies and generate trajectories that are suboptimal or even physically infeasible, especially in long-horizon tasks Talvitie (2014); Asadi et al. (2018); Janner et al. (2019); Voelcker et al. (2022); Chen et al. (2024a).
@@ 2 Planning with Diffusion Models
footnotetext: Codes are available at https://github.com/leekwoon/scots.
@@
10:        $\{\boldsymbol{\tau}_{j}\}_{j=1}^{k}\leftarrow$ TopKNeighbors$(\phi(\text{end}(\boldsymbol{\tau}_{\text{comp}})),\phi(\mathcal{D}),k)$
@@ 3 State-Covering Trajectory Stitching
We introduce State-Covering Trajectory Stitching (SCoTS), a novel reward-free trajectory augmentation framework designed to synthesize an augmented dataset $\mathcal{D}_{\text{aug}}$ from an offline dataset $\mathcal{D}$. The core idea of SCoTS is to iteratively construct long and diverse trajectories by repeatedly stitching short segments guided by latent directional exploration, resulting in significantly improved generalization and extended planning horizons for diffusion planners. SCoTS consists of three stages: (1) learning a temporal distance-preserving embedding for efficient segment retrieval (Section 3.1); (2) iterative trajectory stitching driven by latent directional exploration and novelty-based selection (Section 3.2); and (3) diffusion-based refinement to ensure dynamically consistent transitions (Section 3.3). The overall procedure of SCoTS, including segment search, exploratory selection, and diffusion-based refinement, is illustrated in Figure 2. The detailed algorithm is summarized in Algorithm 1.
@@ 3.1 Temporal Distance-Preserving Embedding
Identifying trajectory segments that are suitable for stitching requires accurately measuring their temporal closeness. However, simply using raw state-space distances can yield temporally incoherent results due to potential dynamic inconsistencies arising from ignoring state reachability. To address this, we employ a temporal distance-preserving embedding $\phi:\mathcal{S}\rightarrow\mathcal{Z}$, which maps raw states to a latent space $\mathcal{Z}$ designed such that the Euclidean distance $\|\phi(\boldsymbol{s})-\phi(\boldsymbol{g})\|_{2}$ approximates the optimal temporal distance $d^{*}(\boldsymbol{s},\boldsymbol{g})$, defined as the minimum number of environment steps required to transition from state $\boldsymbol{s}$ to state $\boldsymbol{g}$. Formally, we parameterize a goal-conditioned value function $V(\boldsymbol{s},\boldsymbol{g})$ following Park et al. (2024b):
$V(\boldsymbol{s},\boldsymbol{g})\coloneqq-||\phi(\boldsymbol{s})-\phi(\boldsymbol{g})||_{2},$ (6)
@@
$\mathcal{L}_{\phi}\coloneqq\mathbb{E}_{(\boldsymbol{s},\boldsymbol{a},\boldsymbol{s}^{\prime},\boldsymbol{g})\sim\mathcal{D}}\left[\ell^{2}_{\xi}(-\mathbbm{1}(\boldsymbol{s}\neq\boldsymbol{g})-\gamma||\bar{\phi}(\boldsymbol{s}^{\prime})-\bar{\phi}(\boldsymbol{g})||_{2}+||\phi(\boldsymbol{s})-\phi(\boldsymbol{g})||_{2})\right],$ (7)
@@ 3.2 Directional and Exploratory Trajectory Stitching
where the distance metric is $\|\phi(\text{end}(\boldsymbol{\tau}_{\text{comp}}))-\phi(\boldsymbol{s}_{1,j})\|_{2}$.
@@ 3.3 Diffusion-based Stitching Refinement
By iteratively repeating segment search, exploratory selection, and this refinement process, we construct a diverse set of augmented trajectories. To generate corresponding action sequences for these trajectories, we train an inverse dynamics model $\boldsymbol{a}_{t}=f_{\psi}(\boldsymbol{s}_{t},\boldsymbol{s}_{t+1})$ on the offline dataset $\mathcal{D}$, which infers the actions that transition between consecutive states. The resulting state-action trajectories are aggregated into the augmented dataset $\mathcal{D}_{\text{aug}}$. This systematic and iterative augmentation approach generates an augmented dataset that broadly covers the state space. Crucially, diffusion planners trained on this augmented data exhibit significantly enhanced trajectory stitching capabilities and improved long-horizon generalization, particularly for tasks requiring extensive trajectory stitching and long-horizon reasoning (Section 4.3).
@@ 4 Experiments
PointMaze Stitch Medium $21$$\pm 9$ ${80}$$\pm 12$ 00 ±$\pm 1$ $74$$\pm 6$ 100100{100}100±$\pm 0$ 100100{100}100±$\pm 0$ 2424{24}24±$\pm 3$ 100100{100}100±$\pm 0$
Large $31$$\pm 2$ ${84}$$\pm 15$ 00 ±$\pm 0$ $13$$\pm 6$ 100100{100}100±$\pm 0$ 100100{100}100±$\pm 0$ 1717{17}17±$\pm 2$ 100100{100}100±$\pm 0$
Giant 00 ±$\pm 0$ ${50}$$\pm 8$ 00 ±$\pm 0$ 00 ±$\pm 0$ 2929{29}29±$\pm 3$ 6868{68}68±$\pm 3$ $\pm 0$ 100100{100}100±$\pm 0$
AntMaze Stitch Medium $29$$\pm 6$ $59$$\pm 7$ $53$$\pm 6$ ${94}$$\pm 1$ 9797{97}97±$\pm 2$ 9696{96}96±$\pm 2$ 7171{71}71±$\pm 1$ 9797{97}97±$\pm 1$
Large $7$$\pm 2$ $18$$\pm 2$ $11$$\pm 2$ ${67}$$\pm 5$ $6${66}66±$\pm 2$ 8686{86}86±$\pm 2$ 3636{36}36±$\pm 2$ 9393{93}93±$\pm 1$
Giant 00 ±$\pm 0$ 00 ±$\pm 0$ 00 ±$\pm 0$ ${2}$$\pm 2$ 2020{20}20±$\pm 1$ 6565{65}65±$\pm 3$ $\pm 0$ 8787{87}87±$\pm 2$
Explore Medium $13$$\pm 2$ $1$$\pm 1$ $3$$\pm 2$ ${37}$$\pm 10$ 9090{90}90±$\pm 2$ 8181{81}81±$\pm 2$ 4242{42}42±$\pm 3$ $9${99}99±$\pm 1$
Large 00 ±$\pm 0$ 00 ±$\pm 0$ 00 ±$\pm 0$ $4$$\pm 5$ 2121{21}21±$\pm 3$ 2727{27}27±$\pm 1$ 1313{13}13±$\pm 2$ 9898{98}98±$\pm 1$
@@
4.1 Datasets and Environments
We evaluate SCoTS on OGBench benchmark Park et al. (2024a), spanning diverse difficulties, environment sizes, agent state dimensions, and training data qualities. Specifically, the benchmark includes three locomotion environments: PointMaze (controlling a 2D point mass) and AntMaze (controlling an 8-DoF quadrupedal Ant). We consider two distinct dataset types, each designed to evaluate specific challenges. The Stitch dataset comprises short, goal-reaching trajectories limited to four cell units, thus requiring the agent to stitch multiple segments (up to 8) for successful inference. In contrast, the Explore dataset assesses learning navigation behaviors from extensive yet low-quality exploratory trajectories, collected by frequently resampling random directions and injecting significant action noise. For each environment, we report the success rate averaged over all evaluation episodes, where an episode is considered successful if the agent reaches sufficiently close to the goal state within a predefined distance threshold. See Appendix A for dataset details.
4.2 Diversity and State Coverage Analysis
To investigate whether SCoTS effectively promotes diverse state-space coverage through trajectory stitching, we evaluate its performance in the PointMaze-Giant-Stitch environment. As illustrated in Figure 4, we visualize the incremental stitching process for different values of the novelty weighting parameter $\beta\in\{0,2,20\}$. We observe that when $\beta=0$, trajectory stitching predominantly follows latent directional guidance, resulting in trajectories with limited coverage but clear directional distinctions. With a moderate setting $\beta=2$, trajectories exhibit a balanced trade-off, achieving substantial state-space coverage with notable diversity. Conversely, at a higher novelty weight $\beta=10$, trajectories broadly cover the state space but lose their distinctiveness, leading to overlapping paths across different latent exploration directions. Based on these results, we use $\beta=2.0$ across all environments in our experiments.
4.3 Diffusion Planning with SCoTS-Augmented Data
@@
We next demonstrate how SCoTS-generated trajectories enhance the ability of diffusion planners to generate feasible, long-horizon plans beyond their training distribution. We compare our approach with offline goal-conditioned reinforcement learning (GCRL) methods including goal-conditioned implicit Q-learning (GCIQL) Kostrikov et al. (2022), Quasimetric RL (QRL) Wang et al. (2023), Contrastive RL (CRL) Eysenbach et al. (2022), and Hierarchical implicit Q-learning (HIQL) Park et al. (2023a). We also include diffusion-based generative planning baselines explicitly designed for long-horizon generalization, such as Generative Skill Chaining (GSC) Mishra et al. (2023) and Compositional Diffuser (CD) Luo et al. (2025).
@@
For our experiments, we adopt a hierarchical diffusion planner (HD) Chen et al. (2024c) that generates plans through a two-level planning process. Specifically, the high-level diffusion model first generates sparse, temporally coarse waypoints, after which a low-level diffusion model fills in the intermediate states between these waypoints, producing a temporally dense trajectory. Initially constrained by limited and short-horizon training data, we augment the original dataset with SCoTS-generated trajectories. After dataset augmentation, we train diffusion planner and employ a value-based low-level controller for action execution, following recent approaches Yoon et al. (2025); Lu et al. (2025). The plans generated by the difussion planner serve as sequences of subgoals for the low-level controller. At each step, the low-level controller executes actions toward a subgoal selected from the generated plan; after a fixed horizon or once the subgoal is reached, it dynamically updates the subgoal by selecting the next state at a specified horizon further along in the plan generated by the diffusion planner. For each dataset, we upsample the original data to 5M samples. Additional implementation details, including hyperparameters and specifics of the low-level controller, are provided in Appendix B.
@@
PointMaze Stitch Medium $21$$\pm 9$ $30$$\pm 3$ $79$$\pm 1$ 00 ±$\pm 1$ 00 ±$\pm 0$ $46$$\pm 2$ $74$$\pm 6$ $77$$\pm 4$ $82$$\pm 4$
Large $31$$\pm 2$ $35$$\pm 4$ $26$$\pm 2$ 00 ±$\pm 0$ 00 ±$\pm 0$ $39$$\pm 2$ $13$$\pm 6$ $16$$\pm 3$ $67$$\pm 1$
Giant 00 ±$\pm 0$ 00 ±$\pm 0$ 00 ±$\pm 0$ 00 ±$\pm 0$ 00 ±$\pm 0$ $18$$\pm 2$ 00 ±$\pm 0$ 00 ±$\pm 0$ $27$$\pm 2$
AntMaze Stitch Medium $29$$\pm 6$ $31$$\pm 3$ $35$$\pm 2$ $53$$\pm 6$ $48$$\pm 3$ $65$$\pm 3$ ${94}$$\pm 1$ $91$$\pm 2$ $94$$\pm 1$
Large $7$$\pm 2$ $3$$\pm 4$ $7$$\pm 1$ $11$$\pm 2$ $12$$\pm 2$ $19$$\pm 1$ ${67}$$\pm 5$ $65$$\pm 3$ $91$$\pm 2$
Giant 00 ±$\pm 0$ 00 ±$\pm 0$ 00 ±$\pm 0$ 00 ±$\pm 0$ 00 ±$\pm 0$ $2$$\pm 1$ ${2}$$\pm 2$ 00 ±$\pm 0$ $55$$\pm 5$
Explore Medium $13$$\pm 2$ $12$$\pm 3$ $18$$\pm 3$ $3$$\pm 2$ $3$$\pm 1$ $15$$\pm 3$ ${37}$$\pm 10$ $45$$\pm 8$ $94$$\pm 1$
Large 00 ±$\pm 0$ 00 ±$\pm 0$ 00 ±$\pm 0$ 00 ±$\pm 0$ $2$$\pm 1$ $19$$\pm 1$ $4$$\pm 5$ $12$$\pm 3$ $77$$\pm 2$

### v3

@@ Abstract
Diffusion-based generative models are emerging as powerful tools for long-horizon planning in reinforcement learning (RL), particularly with offline datasets. However, their performance is fundamentally limited by the quality and diversity of training data. This often restricts their generalization to tasks outside their training distribution or longer planning horizons. To overcome this challenge, we propose State-Covering Trajectory Stitching (SCoTS), a novel reward-free trajectory augmentation method that incrementally stitches together short trajectory segments, systematically generating diverse and extended trajectories. SCoTS first learns a temporal distance-preserving latent representation that captures the underlying temporal structure of the environment, then iteratively stitches trajectory segments guided by directional exploration and novelty to effectively cover and expand this latent space. We demonstrate that SCoTS significantly improves the performance and generalization capabilities of diffusion planners on offline goal-conditioned benchmarks requiring stitching and long-horizon reasoning. Furthermore, augmented trajectories generated by SCoTS significantly improve the performance of widely used offline goal-conditioned RL algorithms across diverse environments. Our code is available at https://github.com/leekwoon/scots/
@@ 1 Introduction
In many real-world applications, agents must plan over hundreds of steps, often receiving sparse or delayed feedback until they reach a distant goal. Perfect knowledge of the environment allows powerful planners like MPC Tassa et al. (2012) and MCTS Silver et al. (2016, 2017) to excel. However, most real-world tasks instead require learning environment dynamics from data. Model-based reinforcement learning (MBRL) Sutton (2018) constructs such world models, offering sample-efficient learning and improved generalization Ha & Schmidhuber (2018); Hafner et al. (2019); Kaiser et al. (2020). However, autoregressive predictions from learned models accumulate small errors into a cascade of inaccuracies. This compounding error can cause planners to exploit model inaccuracies and generate trajectories that are suboptimal or even physically infeasible, especially in long-horizon tasks Talvitie (2014); Asadi et al. (2018); Janner et al. (2019); Voelcker et al. (2022); Chen et al. (2024a).
@@ 2 Planning with Diffusion Models
[…]
@@ 3 State-Covering Trajectory Stitching
We introduce State-Covering Trajectory Stitching (SCoTS), a novel reward-free trajectory augmentation framework designed to synthesize an augmented dataset $\mathcal{D}_{\text{aug}}$ from an offline dataset $\mathcal{D}$. The core idea of SCoTS is to iteratively construct long and diverse trajectories by repeatedly stitching short segments guided by latent directional exploration, resulting in significantly improved generalization and extended planning horizons for diffusion planners. SCoTS consists of three stages: (1) learning a temporal distance-preserving embedding for efficient segment retrieval (Section˜3.1); (2) iterative trajectory stitching driven by latent directional exploration and novelty-based selection (Section˜3.2); and (3) diffusion-based refinement to ensure dynamically consistent transitions (Section˜3.3). The overall procedure of SCoTS, including segment search, exploratory selection, and diffusion-based refinement, is illustrated in Figure 2. The detailed algorithm is summarized in Algorithm 1.
@@ 3.1 Temporal Distance-Preserving Embedding
Identifying trajectory segments that are suitable for stitching requires accurately measuring their temporal closeness. However, simply using raw state-space distances can yield temporally incoherent results due to potential dynamic inconsistencies arising from ignoring state reachability. To address this, we employ a temporal distance-preserving embedding $\phi:\mathcal{S}\rightarrow\mathcal{Z}$, which maps raw states to a latent space $\mathcal{Z}$ designed such that the Euclidean distance $\|\phi(\boldsymbol{s})-\phi(\boldsymbol{g})\|_{2}$ approximates the optimal temporal distance $d^{*}(\boldsymbol{s},\boldsymbol{g})$, defined as the minimum number of environment steps required to transition from state $\boldsymbol{s}$ to state $\boldsymbol{g}$. Formally, we parameterize a goal-conditioned value function $V(\boldsymbol{s},\boldsymbol{g})$ following Park et al. (2024a):
[…]
@@
[…]
@@
Remark.
We note that the learned latent space is not a perfect metric representation of the MDP Park et al. (2024a). However, the reliability of SCoTS is grounded in its design, which does not require a globally accurate temporal distance metric. Instead, as we detail next, our framework leverages the latent distance for the more tractable local problem of retrieving promising and reachable candidate segments at each step of the iterative stitching process. This local approach makes the overall framework robust to the inherent imperfections of the embedding.
@@ 3.2 Directional and Exploratory Trajectory Stitching
where $\phi(\mathcal{D})$ is a concise representation for the set of latent embeddings of the initial states of all trajectories within the dataset $\mathcal{D}$, and the distance metric is $\|\phi(\text{end}(\boldsymbol{\tau}_{\text{comp}}))-\phi(\boldsymbol{s}_{1,j})\|_{2}$.
@@ 3.3 Diffusion-based Stitching Refinement
By iteratively repeating segment search, exploratory selection, and this refinement process, we construct a diverse set of augmented trajectories. To generate corresponding action sequences for these trajectories, we train an inverse dynamics model $\boldsymbol{a}_{t}=f_{\psi}(\boldsymbol{s}_{t},\boldsymbol{s}_{t+1})$ on the offline dataset $\mathcal{D}$, which infers the actions that transition between consecutive states. The resulting state-action trajectories are aggregated into the augmented dataset $\mathcal{D}_{\text{aug}}$. This systematic and iterative augmentation approach generates an augmented dataset that broadly covers the state space. Crucially, diffusion planners trained on this augmented data exhibit significantly enhanced trajectory stitching capabilities and improved long-horizon generalization, particularly for tasks requiring extensive trajectory stitching and long-horizon reasoning (Section˜4.3).
@@ 4 Experiments
4.1 Datasets and Environments
We evaluate SCoTS on OGBench benchmark Park et al. (2025), spanning diverse difficulties, environment sizes, agent state dimensions, and training data qualities. Specifically, the benchmark includes three locomotion environments: PointMaze (controlling a 2D point mass) and AntMaze (controlling an 8-DoF quadrupedal Ant). We consider two distinct dataset types, each designed to evaluate specific challenges. The Stitch dataset comprises short, goal-reaching trajectories limited to four cell units, thus requiring the agent to stitch multiple segments (up to 8) for successful inference. In contrast, the Explore dataset assesses learning navigation behaviors from extensive yet low-quality exploratory trajectories, collected by frequently resampling random directions and injecting significant action noise. For each environment, we report the success rate averaged over all evaluation episodes, where an episode is considered successful if the agent reaches sufficiently close to the goal state within a predefined distance threshold. See Appendix A for dataset details.
4.2 Diversity and State Coverage Analysis
To investigate whether SCoTS effectively promotes diverse state-space coverage through trajectory stitching, we evaluate its performance in the PointMaze-Giant-Stitch environment. As illustrated in Figure 3, we visualize the incremental stitching process for different values of the novelty weighting parameter $\beta\in\{0,2,20\}$. We observe that when $\beta=0$, trajectory stitching predominantly follows latent directional guidance, resulting in trajectories with limited coverage but clear directional distinctions. With a moderate setting $\beta=2$, trajectories exhibit a balanced trade-off, achieving substantial state-space coverage with notable diversity. Conversely, at a higher novelty weight $\beta=20$, trajectories broadly cover the state space but lose their distinctiveness, leading to overlapping paths across different latent exploration directions. Based on these results, we use $\beta=2$ across all environments in our experiments.
4.3 Diffusion Planning with SCoTS-Augmented Data
@@
[…]
@@
We next demonstrate how SCoTS-generated trajectories enhance the ability of diffusion planners to generate feasible, long-horizon plans beyond their training distribution. We compare our approach with offline goal-conditioned reinforcement learning (GCRL) methods including goal-conditioned implicit Q-learning (GCIQL) Kostrikov et al. (2022), Quasimetric RL (QRL) Wang et al. (2023), Contrastive RL (CRL) Eysenbach et al. (2022), and Hierarchical implicit Q-learning (HIQL) Park et al. (2023). We also include diffusion-based generative planning baselines explicitly designed for long-horizon generalization, such as Generative Skill Chaining (GSC) Mishra et al. (2023) and Compositional Diffuser (CD) Luo et al. (2025).
@@
For our experiments, we adopt a hierarchical diffusion planner (HD) Chen et al. (2024c) that generates plans through a two-level planning process. Specifically, the high-level diffusion model first generates sparse, temporally coarse waypoints, after which a low-level diffusion model fills in the intermediate states between these waypoints, producing a temporally dense trajectory. Initially constrained by limited and short-horizon training data, we augment the original dataset with SCoTS-generated trajectories. After dataset augmentation, we train diffusion planner and employ a value-based low-level controller for action execution, following recent approaches Yoon et al. (2025a); Lu et al. (2025). The plans generated by the diffusion planner serve as sequences of subgoals for the low-level controller. At each step, the low-level controller executes actions toward a subgoal selected from the generated plan; after a fixed horizon or once the subgoal is reached, it dynamically updates the subgoal by selecting the next state at a specified horizon further along in the plan generated by the diffusion planner. For each dataset, we upsample the original data to 5M samples. Additional implementation details, including hyperparameters and specifics of the low-level controller, are provided in Appendix B.
@@
PointMaze Stitch Medium 2121 $\pm 9$ $30$ $\pm 3$ $79$ $\pm 1$ 0 $\pm 1$ 0 $\pm 0$ $46$ $\pm 2$ $74$ $\pm 6$ $77$ $\pm 4$ $82$ $\pm 4$
Large $31$ $\pm 2$ $35$ $\pm 4$ $26$ $\pm 2$ 0 $\pm 0$ 0 $\pm 0$ $39$ $\pm 2$ $13$ $\pm 6$ $16$ $\pm 3$ $67$ $\pm 1$
Giant 0 $\pm 0$ 0 $\pm 0$ 0 $\pm 0$ 0 $\pm 0$ 0 $\pm 0$ $18$ $\pm 2$ 0 $\pm 0$ 0 $\pm 0$ $27$ $\pm 2$
AntMaze Stitch Medium $29$ $\pm 6$ $31$ $\pm 3$ $35$ $\pm 2$ $53$ $\pm 6$ $48$ $\pm 3$ $65$ $\pm 3$ 94{94} $\pm 1$ $91$ $\pm 2$ $94$ $\pm 1$
Large 77 $\pm 2$ 33 $\pm 4$ 77 $\pm 1$ $11$ $\pm 2$ $12$ $\pm 2$ 1919 $\pm 1$ 67{67} $\pm 5$ $65$ $\pm 3$ $91$ $\pm 2$
Giant 0 $\pm 0$ 0 $\pm 0$ 0 $\pm 0$ 0 $\pm 0$ 0 $\pm 0$ 22 $\pm 1$ 2{2} $\pm 2$ 0 $\pm 0$ $55$ $\pm 5$
Explore Medium $13$ $\pm 2$ $12$ $\pm 3$ $18$ $\pm 3$ 33 $\pm 2$ 33 $\pm 1$ $15$ $\pm 3$ 37{37} $\pm 10$ $45$ $\pm 8$ $94$ $\pm 1$
Large 0 $\pm 0$ 0 $\pm 0$ 0 $\pm 0$ 0 $\pm 0$ 22 $\pm 1$ 1919 $\pm 1$ 44 $\pm 5$ $12$ $\pm 3$ $77$ $\pm 2$